OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6

OpenStudy (anonymous):

multiply second eqn by 2 and add both d eqn...ull get value of x.... den put it into any of eqn...ull get value of y..

OpenStudy (anonymous):

-3/2x-2y=2 2y=-3/2x-2 y=(-3/2x-2)/2

OpenStudy (anonymous):

multiplaying 2nd eqn by 2 ull get -4x+6y=12 if u add 1st eqn and dis eqn... 3x - 6y = 12 + -4x+6y=12 y term wil get cancel s dey hv opposite signs....and ull get -x=24 that is x=-24 put dis value of x in any of eqn u wil get value of y....

OpenStudy (anonymous):